OpenStudy (anonymous):

square root of 45x^3y

OpenStudy (anonymous):

\[45=9\times 5\] so \(\sqrt{45}=\sqrt{9\times 5}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\) is a start

OpenStudy (anonymous):

also \(\sqrt{x^3}=\sqrt{x^2}\times \sqrt{x}=x\sqrt{x}\)

OpenStudy (anonymous):

so a \(3\) comes out of the radical, a \(x\) comes outside, and the \(5xy\) stays inside, giving \[3x\sqrt{5xy}\]

OpenStudy (anonymous):

can u help with the whole problem? its √45x^3y - √20xy/ √2xy